• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 07:38:25 on June16,2010

    Entry number 326887

    keyword=owl shift summary

    Production on Pb/D tgt 1

    2-HRS running until about 1:30 am. Then 1-HRS (L).

    0:46 ... IWHP to IN, slug 36 started

    02:00 Access to try to fix R-HRS dipole. Failed to fix it.

    03:00 Resume production with L-HRS only.

    06:20 ACC down. Vacuum problems in South Linac.

    07:30 Howard Smith (tech) calls in. We decide to go
    to controlled access to fix the R-HRS dipole.

    Parity runs
    4587 last run of slug 35
    4588 - 4594 with beam (mostly)
    4595 pedestal

    Compton runs (long runs)
    23137 - 23139

    Note: when the beam returns don't forget to

    -- turn off heater and turn on cooler
    -- start a new Compton DAQ run (we stopped because of no beam)